# Hyperbolic Fuzzy $C$-Means with Adaptive Weight-based Filtering for Clustering in Non-Euclidean Spaces

### v1

@@ 2.3 Hyperbolic space
The parameter $m\in[2,\infty)$ in [34] is the fuzziness parameter. According to [35], the value of $m$ is usually taken as 2. The optimization problem in Equation 12 can be solved using Lagrange multipliers [36]. Using the Alternative Minimization method, the update formula for the centroid is given by :
@@ 3.2 HSFC
Consider the hyperbolic smoothing function $\phi_{(}y,\tau)=\frac{y+\sqrt{y^{2}+\tau^{2}}}{2}$, for all $y\in\mathbb{R},\tau\geq 0$ and the function: $\theta\left(\mathbf{x}_{i},\mathbf{v}_{j},\gamma\right)=$ $\sqrt{\sum_{k=1}^{p}\left(x_{ik}-v_{jk}\right)^{2}+\gamma^{2}}$, for $\gamma>0$. Hence, the Minimization Problem of Hyperbolic Smoothing Clustering Method (HSFC) [37] is transformed into:
@@
3.3 P_SFCM
Assume a data matrix is indicated as $\mathbf{X}\in\mathbb{R}^{n\times p}$, where $n$ is the number of data points and $p$ indicates the dimensionality of the data. $\mathbf{W}=(w_{ij})$ is the $(n\times c)$ membership matrix. Hence, the Minimization Problem of P_SFCM is given as :
$\min_{\mathbf{P},\mathbf{W},\mathbf{V}}\sum_{i=1}^{n}f_{i}\sum_{j=1}^{c}\|\mathbf{x}_{i}\mathbf{P}^{T}-\mathbf{\tilde{v}}_{j}\|_{2}^{2}(w_{ij})^{m}-\lambda Tr(\mathbf{P}\mathbf{S}_{t}\mathbf{P}^{T}),$
[…]
where, $\mathbf{P}\in\mathbb{R}^{q\times p}$ is the projection matrix, such that $\mathbf{X}\mathbf{P}^{T}\in\mathbb{R}^{n\times q},q\leq p$, sum of each row of $\mathbf{W}$ is 1, $\mathbf{\tilde{v}}_{j}$ denotes the cluster centriods, $f_{i}$ is the element of the matrix $\mathbf{F}$ s.t. $f_{i}\in\{0,1\}.$
4 HypeFCM Method and its algorithm.
In this section, we will present our proposed clustering algorithm, HypeFCM with intricate details.
@@ 3.3 P_SFCM
Input: Data (Xn×p), $k$ = filtration value, $c$ = number of clusters, $m$ = fuzziness parameter, $T$ = maximum number of iterations, $\varepsilon$ = tolerance value.
Output: The clusters $Y_{1},Y_{2},\cdots Y_{c}.$
1: Embed the original dataset $\mathbf{X}$ into the Poincaré Disc of radius $\frac{1}{\sqrt{\alpha}}$. Obtain $\mathbf{X^{\prime}}=\{\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime},\cdots,\mathbf{x}_{n}^{\prime}\}\in\mathbb{D}_{\alpha}^{p}.$
@@
3: Obtain the centroids $\{\mathbf{v}^{(t)}_{j}\}_{j=1}^{c}$ using the Riemannian log-exp map:
$\bar{v}_{j}=\frac{\sum_{i=1}^{n}(w_{ij}^{(t)})^{m}\cdot\log_{\mathbf{v}_{j}^{(t-1)}}(\mathbf{x}_{i}^{\prime})}{\sum_{i=1}^{n}(w_{ij}^{(t)})^{m}}$
$\mathbf{v}_{j}^{(t)}=\exp_{\mathbf{v}_{j}^{(t-1)}}(\bar{v}_{j})$
where, $\log$ and $\exp$ denote the logarithmic and exponential maps on the Poincaré ball respectively.
4: Obtain the distance matrix, U(t) by considering distances from the centroids,
$\mathbf{U^{(t)}_{n\times c}}:=u^{(t)}_{ij}:=d_{H}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}^{(t)}_{j}))_{i\in[n],j\in[c]}.$
using the squared hyperbolic distance between data points and centroids.
5: Apply filtration based on the distances of data points from each of the centroids and the rest of the entries of U(t), are assigned to zero to obtain $\mathbf{U}^{\prime(t)}$.
6: Update the weights using:
$w^{(t+1)}_{ij}=\frac{(u^{\prime(t)}_{ij})^{-1/(m-1)}}{\sum_{j}(u^{\prime(t)}_{ij})^{-1/(m-1)}}.$
7:for $t$ from $1$ to $T$ do
       Update $\mathbf{W}^{(t)}$ by Step ($6$)
      Update centroids $\mathbf{V}^{(t)}:=\{\mathbf{v}^{(t)}_{j}\}_{j=1}^{c}$ by Step ($3$)
      if $||\mathbf{W}^{(t+1)}-\mathbf{W}^{(t)}||\leq\varepsilon$ then
            break
       end if
end for
8: Assign labels by selecting the index of the maximum membership value in each row of the finally updated weight matrix.
@@
$\mathbf{v}_{j}^{(t+1)}=\exp_{\mathbf{v}_{j}^{(t)}}\left(\frac{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}\log_{\mathbf{v}_{j}^{(t)}}(\mathbf{x}^{\prime}_{i})}{\sum_{i=1}^{n}(w_{ij}^{(t+1)})^{m}}\right).$ (17)
When limit $\alpha$ tends to 00, this update formula 17 becomes:
@@
which is the update formula for the centroid in FCM in Euclidean space using the equations 10, 11.
@@
$\mathbf{U^{(t)}_{n\times c}}:=u^{(t)}_{ij}:=d_{H}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}^{(t)}_{j})_{i\in[n],j\in[c]},$
where, $d_{H}$ is the hyperbolic distance mentioned in Equation 7. In each row of the matrix $\mathbf{U}$, the nearest $k$ data points from the respective centroid are kept as it is and the remaining entries are assigned to 00, that is, given the $t^{th}$ iteration of the centroids, only those data points that are the least distant from the respective centroid are kept and the rest are zero.
In essence, we define a relation $\mathbf{R}$ on the Cartesian product of the set of the centroids and the set of the datapoints, where a datapoint $i$ and a centroid $j$ are related such that,
@@
$u^{\prime}_{ij}=\Bigl{(}d_{H}^{2}(\mathbf{x}_{i}^{\prime}-\mathbf{v}_{j})\mathds{1}\{(i,j)\in\mathbf{R}\}\Bigr{)}\quad\forall i\in[n],\forall j\in[c],$
where ${.}\mathds{1\{.\}}$ denotes the indicator function and $\mathbf{R}$ being the relation as described above. Then the expression for the updation of the weight matrix $\mathbf{W}$ is given as:
@@
Figure 1: The t-SNE visualization of real datasets (a) Glass, (b) Ecoli, and (c) Wine respectively, for FCM and HypeFCM clustering methods.
@@
Iteration Step. Performing steps $3$-$6$ for each iteration in Algorithm 1 requires time complexity of $\mathcal{O}(ncp+nc\log k)$. Thus, the total time complexity for $T$ iterations becomes $\mathcal{O}(T(ncp+nc\log k))$.
@@
Therefore, the overall time complexity for our HypeFCM algorithm is $\mathcal{O}(T(ncp+nc\log k))$, equivalent to $\mathcal{O}(Tncp)$ since $k\ll n$, which is comparable to the overall time complexities of the FCM and PCM algorithms.
@@
Table I: Comparison of Clustering Performance across multiple methods, $K$-means, FCM, PCM, P_SFCM, MiniBatchKmeans, HSFC, with our proposed HypeFCM ON $6$ Synthetic datasets, Presented as Means with Standard Deviations.
@@
Datasets Metric FCM PCM P_SFCM $K$-means MiniBatchKmeans HSFC HypeFCM (Ours)
Cure-t1-2000n-2D ARI 0.471 $\pm$ 0.014 0.079 $\pm$ 0.065 0.486 $\pm$ 0.042 0.491 $\pm$ 0.021 0.505 $\pm$ 0.019 0.493 $\pm$ 0.022 0.723 $\pm$ 0.030
NMI 0.574 $\pm$ 0.012 0.074 $\pm$ 0.059 0.640 $\pm$ 0.038 0.747 $\pm$ 0.018 0.774 $\pm$ 0.016 0.762 $\pm$ 0.019 0.781 $\pm$ 0.039
Cure-t2-4k ARI 0.428 $\pm$ 0.018 0.068 $\pm$ 0.040 0.447 $\pm$ 0.056 0.438 $\pm$ 0.035 0.532 $\pm$ 0.023 0.503 $\pm$ 0.037 0.581 $\pm$ 0.007
NMI 0.607 $\pm$ 0.017 0.072 $\pm$ 0.044 0.608 $\pm$ 0.052 0.666 $\pm$ 0.031 0.649 $\pm$ 0.020 0.617 $\pm$ 0.033 0.676 $\pm$ 0.006
Donutcurves ARI 0.358 $\pm$ 0.021 0.070 $\pm$ 0.035 0.498 $\pm$ 0.037 0.420 $\pm$ 0.032 0.423 $\pm$ 0.031 0.461 $\pm$ 0.019 0.501 $\pm$ 0.006
NMI 0.514 $\pm$ 0.019 0.075 $\pm$ 0.044 0.663 $\pm$ 0.043 0.576 $\pm$ 0.028 0.581 $\pm$ 0.027 0.634 $\pm$ 0.025 0.667 $\pm$ 0.005
Smile1 ARI 0.542 $\pm$ 0.015 0.142 $\pm$ 0.064 0.551 $\pm$ 0.041 0.527 $\pm$ 0.026 0.551 $\pm$ 0.024 0.548 $\pm$ 0.027 0.594 $\pm$ 0.006
NMI 0.606 $\pm$ 0.016 0.325 $\pm$ 0.059 0.612 $\pm$ 0.037 0.559 $\pm$ 0.022 0.607 $\pm$ 0.021 0.607 $\pm$ 0.023 0.701 $\pm$ 0.005
Disk-4000n ARI 0.001 $\pm$ 0.002 0.016 $\pm$ 0.007 0.001 $\pm$ 0.019 -0.002 $\pm$ 0.018 -0.001 $\pm$ 0.025 0.010 $\pm$ 0.025 0.239 $\pm$ 0.033
NMI 0.002 $\pm$ 0.004 0.017 $\pm$ 0.005 0.002 $\pm$ 0.015 0.001 $\pm$ 0.012 0.002 $\pm$ 0.022 0.101 $\pm$ 0.021 0.293 $\pm$ 0.043
3MC ARI 0.584 $\pm$ 0.013 0.142 $\pm$ 0.068 0.553 $\pm$ 0.035 0.508 $\pm$ 0.019 0.546 $\pm$ 0.018 0.498 $\pm$ 0.020 0.586 $\pm$ 0.008
NMI 0.650 $\pm$ 0.012 0.325 $\pm$ 0.062 0.612 $\pm$ 0.033 0.608 $\pm$ 0.017 0.628 $\pm$ 0.016 0.531 $\pm$ 0.018 0.653 $\pm$ 0.009
    Average Rank 4.75 6.33 3.83 4.17 3.25 3.67 1.08
@@
6 Convergence Analysis.
We will introduce the objective function of HypeFCM. Let $\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots\mathbf{x}_{n}\}$ be a sample of $n$ observations in $\mathbb{D}_{\alpha}^{p}$. $\mathbf{V}$ is an $(c\times p)$ matrix that represents the centroids of the $c$ clusters in $\mathbb{D}_{\alpha}^{p}$ and $\mathbf{W}$ is an $(n\times c)$ membership matrix with elements $w_{ij}\in[0,1]$ such that the following least squares criterion is minimized. In this section, we offer insights into the convergence analysis of HypeFCM.
$Minimize\hskip 5.0ptJ_{m}(\mathbf{X},\mathbf{W},\mathbf{V})=\sum_{i=1}^{n}\sum_{j=1}^{c}(w_{ij})^{m}d_{H}^{2}(\mathbf{x}_{i},\mathbf{v}_{j}),$

### v2

@@ 3.1 Fuzzy $C$-means (FCM)
The parameter $m\in[2,\infty)$ in [37] is the fuzziness parameter. According to [38], the value of $m$ is usually taken as 2. Using the Alternative Minimization method, the update formula for the centroid is given by :
@@ 3.2 HSFC
Consider the hyperbolic smoothing function $\phi_{(}y,\tau)=\frac{y+\sqrt{y^{2}+\tau^{2}}}{2}$, for all $y\in\mathbb{R},\tau\geq 0$ and the function:
$\theta\left(\mathbf{x}_{i},\mathbf{v}_{j},\gamma\right)=$ $\sqrt{\sum_{k=1}^{p}\left(x_{ik}-v_{jk}\right)^{2}+\gamma^{2}}$, for $\gamma>0$. Hence, the Minimization Problem of Hyperbolic Smoothing Clustering Method (HSFC) [39] is transformed into:
@@
3.3 FCSR
The minimization function of the Fuzzy Clustering Guided by Spectral Rotation and Scaling (FCSR) [40] model is defined as:
$J_{\text{FCSR}}(W,V,R,\Phi)$ $=\sum_{i=1}^{n}\sum_{k=1}^{c}\|\mathbf{x}_{i}-\mathbf{v}_{k}\|_{2}^{2}\,w_{ik}^{2}$ (13)
$+\lambda\|W-\Phi FR\|_{F}^{2}$
s.t. $W\geq 0,\quad W\mathbf{1}=\mathbf{1},$
$R^{\top}R=I,$
$\Phi=\text{diag}(\phi_{11},\phi_{22},\ldots,\phi_{nn}),\quad\phi_{ii}\geq 0$
$\quad(i=1,2,\ldots,n),$
where, $F\in\mathbb{R}^{n\times c}$ is the spectral embedding matrix, $R\in\mathbb{R}^{c\times c}$ is the rotation matrix, $W\in\mathbb{R}^{d\times d^{\prime}}$ is the projection matrix to a $d^{\prime}$-dimensional subspace, $\Phi\in\mathbb{R}^{n\times n}$ is the diagonal scaling matrix.
3.4 EFKM
The objective function of the proposed EFKM [41] (Fuzzy $K$-Means Clustering with Discriminative Embedding) is defined as:
$\min_{\begin{subarray}{c}P^{\top}P=I,\,W\mathbf{1}_{c}=\mathbf{1}_{n},\,W\geq 0,\\ M=[\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{c}]\end{subarray}}\frac{\displaystyle\sum_{i=1}^{n}\sum_{j=1}^{c}w_{ij}^{m}\left\|P^{\top}\mathbf{x}_{i}-\textbf{v}_{j}\right\|_{2}^{2}}{\operatorname{Tr}(P^{\top}XX^{\top}P)}$ (14)
where, $P\in\mathbb{R}^{d\times d^{\prime}}$ is the projection matrix to a $d^{\prime}$-dimensional subspace, $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix, The constraints $W\geq 0$, $W\mathbf{1}_{c}=\mathbf{1}_{n}$ ensure that each row of $W$ sums to 1.
Figure 1: The t-SNE visualization of real datasets (a) Glass, (b) Wisconsin Breast Cancer, respectively, for FCM and HypeFCM clustering methods.
4 HypeFCM: our proposed method
In this section, we will present our proposed clustering algorithm, HypeFCM, with intricate details.
@@ 4 HypeFCM: our proposed method
$\bar{v}_{j}=\frac{\sum_{i=1}^{n}(w_{ij}^{(t)})^{m}\cdot\log_{\mathbf{v}_{j}^{(t-1)}}(\mathbf{x}_{i}^{\prime})}{\sum_{i=1}^{n}(w_{ij}^{(t)})^{m}}$ (15)
$\mathbf{v}_{j}^{(t)}=\exp_{\mathbf{v}_{j}^{(t-1)}}(\bar{v}_{j})$ (16)
When limit $\alpha$ tends to 0, this update formula 16 becomes:
@@
which conforms to the update formula for the centroid in FCM in Euclidean space, using the Equations 7 and 8.
@@
$\mathbf{U^{(t)}_{n\times c}}:=u^{(t)}_{ij}:=d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}^{(t)}_{j})_{i\in[n],j\in[c]},$
where, $d_{hyp}$ is the hyperbolic distance mentioned in Equation 4. In each row of the matrix $\mathbf{U}$, the nearest $k$ data points from the respective centroid are kept as it is. The remaining entries are assigned to 0, that is, given the $t^{th}$ iteration of the centroids, only those data points that are the least distant from the respective centroid are kept and the rest are zero.
In essence, we define a relation $\mathbf{R}$ on the Cartesian product of the set of the centroids and the set of the datapoints, where the $i^{th}$ datapoint and the $j^{th}$ centroid are related such that,
@@
$u^{\prime}_{ij}=\Bigl{(}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime}-\mathbf{v}_{j})\mathds{1}\{(i,j)\in\mathbf{R}\}\Bigr{)}\quad\forall i\in[n],\forall j\in[c],$
where $\mathds{1\{.\}}$ denotes the indicator function and $\mathbf{R}$ being the relation as described above. Then the expression for the updation of the weight matrix $\mathbf{W}$ is given as:
@@
Iteration Step. Performing steps 33-66 for each iteration in Algorithm 1 requires time complexity of $\mathcal{O}(ncp+nc\log k)$. Thus, the total time complexity for $T$ iterations becomes $\mathcal{O}(T(ncp+nc\log k))$.
@@
Therefore, the overall time complexity for our HypeFCM algorithm is $\mathcal{O}(T(ncp+nc\log k))$, equivalent to $\mathcal{O}(Tncp)$, since $k\ll n$, which is equivalent to the overall time complexity of the FCM algorithm in Euclidean space.
@@
Table II: Comparison of Clustering Performance across multiple methods, FCM, PCM, P_SFCM, $K$-means, MiniBatchKmeans, EFKM, UFCER, FCSR, IFKMHC, HSFC, FCPFM, with our proposed HypeFCM on $3$ synthetic datasets along with 77 Real-World datasets, presented as Means with Standard Deviations. The best and second-best results are highlighted in boldface and underlined, respectively.
@@
Datasets Metric FCM PCM P_SFCM $K$-means MiniBatchKmeans EFKM UFCER FCSR IFKMHC HSFC FCPFM HypeFCM (Ours)
[…]
Iris ARI 0.810 $\pm$ 0.012 0.468 $\pm$ 0.071 0.680 $\pm$ 0.058 0.742 $\pm$ 0.023 0.800 $\pm$ 0.015 0.517 $\pm$ 0.015 0.508 $\pm$ 0.015 0.759 $\pm$ 0.015 0.604 $\pm$ 0.015 0.742 $\pm$ 0.021 0.667 $\pm$ 0.046 0.812 $\pm$ 0.015
NMI 0.815 $\pm$ 0.015 0.548 $\pm$ 0.065 0.725 $\pm$ 0.051 0.756 $\pm$ 0.027 0.788 $\pm$ 0.019 0.574 $\pm$ 0.015 0.593 $\pm$ 0.015 0.728 $\pm$ 0.015 0.616 $\pm$ 0.015 0.756 $\pm$ 0.025 0.678 $\pm$ 0.055 0.817 $\pm$ 0.014
Glass ARI 0.228 $\pm$ 0.045 0.005 $\pm$ 0.010 0.129 $\pm$ 0.022 0.270 $\pm$ 0.041 0.210 $\pm$ 0.047 0.448 $\pm$ 0.015 0.168 $\pm$ 0.015 0.158 $\pm$ 0.015 0.113 $\pm$ 0.015 0.264 $\pm$ 0.052 0.176 $\pm$ 0.065 0.598 $\pm$ 0.007
NMI 0.334 $\pm$ 0.038 0.035 $\pm$ 0.016 0.156 $\pm$ 0.026 0.428 $\pm$ 0.036 0.356 $\pm$ 0.042 0.491 $\pm$ 0.015 0.297 $\pm$ 0.015 0.311 $\pm$ 0.015 0.151 $\pm$ 0.015 0.398 $\pm$ 0.047 0.302 $\pm$ 0.028 0.609 $\pm$ 0.016
Ecoli ARI 0.378 $\pm$ 0.041 0.390 $\pm$ 0.047 0.293 $\pm$ 0.028 0.384 $\pm$ 0.039 0.390 $\pm$ 0.038 0.354 $\pm$ 0.015 0.465 $\pm$ 0.015 0.286 $\pm$ 0.015 0.351 $\pm$ 0.015 0.424 $\pm$ 0.057 0.384 $\pm$ 0.045 0.518 $\pm$ 0.006
NMI 0.472 $\pm$ 0.034 0.349 $\pm$ 0.061 0.324 $\pm$ 0.034 0.534 $\pm$ 0.035 0.525 $\pm$ 0.031 0.456 $\pm$ 0.015 0.487 $\pm$ 0.015 0.504 $\pm$ 0.015 0.466 $\pm$ 0.015 0.547 $\pm$ 0.039 0.576 $\pm$ 0.025 0.591 $\pm$ 0.010
Wine ARI 0.366 $\pm$ 0.032 0.042 $\pm$ 0.019 0.287 $\pm$ 0.015 0.352 $\pm$ 0.042 0.365 $\pm$ 0.036 0.408 $\pm$ 0.015 0.351 $\pm$ 0.015 0.391 $\pm$ 0.015 0.295 $\pm$ 0.015 0.375 $\pm$ 0.038 0.421 $\pm$ 0.012 0.412 $\pm$ 0.011
NMI 0.425 $\pm$ 0.036 0.089 $\pm$ 0.024 0.320 $\pm$ 0.019 0.423 $\pm$ 0.037 0.430 $\pm$ 0.032 0.498 $\pm$ 0.015 0.365 $\pm$ 0.015 0.428 $\pm$ 0.015 0.351 $\pm$ 0.015 0.428 $\pm$ 0.034 0.481 $\pm$ 0.016 0.457 $\pm$ 0.012
Zoo ARI 0.527 $\pm$ 0.025 0.302 $\pm$ 0.054 0.157 $\pm$ 0.067 0.714 $\pm$ 0.022 0.680 $\pm$ 0.044 0.343 $\pm$ 0.015 0.675 $\pm$ 0.015 0.591 $\pm$ 0.015 0.611 $\pm$ 0.015 0.447 $\pm$ 0.062 0.671 $\pm$ 0.035 0.726 $\pm$ 0.018
NMI 0.543 $\pm$ 0.029 0.457 $\pm$ 0.049 0.376 $\pm$ 0.058 0.782 $\pm$ 0.025 0.747 $\pm$ 0.039 0.427 $\pm$ 0.015 0.711 $\pm$ 0.015 0.615 $\pm$ 0.015 0.664 $\pm$ 0.015 0.482 $\pm$ 0.058 0.723 $\pm$ 0.045 0.788 $\pm$ 0.025
Flights (5k) ARI 0.024 $\pm$ 0.008 -0.003 $\pm$ 0.006 0.014 $\pm$ 0.006 0.013 $\pm$ 0.007 0.015 $\pm$ 0.010 0.035 $\pm$ 0.012 0.002 $\pm$ 0.004 0.021 $\pm$ 0.012 0.013 $\pm$ 0.007 0.029 $\pm$ 0.006 0.007 $\pm$ 0.015 0.042 $\pm$ 0.014
NMI 0.017 $\pm$ 0.006 0.031 $\pm$ 0.012 0.011 $\pm$ 0.006 0.026 $\pm$ 0.008 0.034 $\pm$ 0.010 0.048 $\pm$ 0.012 0.014 $\pm$ 0.006 0.024 $\pm$ 0.009 0.037 $\pm$ 0.015 0.042 $\pm$ 0.015 0.021 $\pm$ 0.011 0.050 $\pm$ 0.016
Phishing (5k) ARI 0.000 $\pm$ 0.003 0.000 $\pm$ 0.015 0.060 $\pm$ 0.046 0.001 $\pm$ 0.001 0.001 $\pm$ 0.002 0.169 $\pm$ 0.015 0.013 $\pm$ 0.009 0.231 $\pm$ 0.015 0.001 $\pm$ 0.002 0.002 $\pm$ 0.001 0.307 $\pm$ 0.015 0.316 $\pm$ 0.021
NMI 0.010 $\pm$ 0.005 0.003 $\pm$ 0.008 0.069 $\pm$ 0.031 0.001 $\pm$ 0.007 0.003 $\pm$ 0.001 0.137 $\pm$ 0.015 0.035 $\pm$ 0.015 0.179 $\pm$ 0.015 0.012 $\pm$ 0.005 0.005 $\pm$ 0.002 0.274 $\pm$ 0.012 0.281 $\pm$ 0.018
Average Rank 6.975 10.687 8.375 6.112 5.375 4.262 7.687 6.562 8.137 5.812 4.937 1.457
@@
6 Convergence Analysis of HypeFCM
We will introduce the objective function of HypeFCM. Let $\mathbf{X^{\prime}}=\{\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime},\ldots\mathbf{x}_{n}^{\prime}\}$ be a sample of $n$ observations in $\mathbb{D}_{\alpha}^{p}$. $\mathbf{V}$ is an $(c\times p)$ matrix that represents the centroids of the $c$ clusters in $\mathbb{D}_{\alpha}^{p}$ and $\mathbf{W}$ is an $(n\times c)$ membership matrix with elements $w_{ij}\in[0,1]$ such that the following least squares criterion is minimized. In this section, we offer insights into the convergence analysis of HypeFCM.
$Minimize\hskip 5.0ptJ_{m}(\mathbf{X^{\prime}},\mathbf{W},\mathbf{V})=\sum_{i=1}^{n}\sum_{j=1}^{c}(w_{ij})^{m}d_{hyp}^{2}(\mathbf{x}_{i}^{\prime},\mathbf{v}_{j}),$